# Deep Linear Hawkes Processes

### v1

@@ B.2 Discretization and Zero Order Hold
The linear recurrence is defined in continuous-time. This mirrors the (M)TPP setting, where event times are not on a fixed intervals. We use the zero-order hold (ZOH) discretization method, to convert the continuous-time linear recurrence into a sequence of closed-form updates, given the integration times, that can also be efficiently computed. We refer the reader to Iserles (2009) for a comprehensive introduction to the ZOH transform.
The main assumption of the ZOH discretization is that the input signal is held constant over the time period being integrated. Under this assumption, it is possible to solve for the dynamics and input matrices that yield the correct state at the end of the integration period. For the LLH dynamics in Eq. 10, when no events occur in $(t,t^{\prime})$, this becomes
@@
The ZOH does not affect the output or passthrough matrices $\mathbf{C}$ and $\mathbf{D}$. To compute the matrices $\overline{\mathbf{A}}$ and $\overline{\mathbf{AB}}$ however requires computing a matrix exponential and a matrix inverse. However, Smith et al. (2022) avoid this by diagonalizing the system (also avoiding a dense matrix-matrix multiplication in the parallel scan). The diagonalized dynamics and input matrices are denoted $\bm{\Lambda}$ (a diagonal matrix) and $\bm{\Lambda}\tilde{\mathbf{B}}$ respectively. In this case, Eq. 19 reduces to
@@
[…]
@@
To contextualize, suppose an event occurs at time $t$, Eq. 22 allows us to exactly (under the constant-input assumption) efficiently evaluate the linear recurrence at subsequent times $t^{\prime}$. We use this extensively in the DLHP to efficiently evaluate the recurrence (and hence the intensity) at the irregularly-spaced event times and times used to compute the integral term.
It should be noted the discretization was done to compute a left-limit $\mathbf{x}_{t^{\prime}-}$ from a previous right-limit $\mathbf{x}_{t}$. Should an event not occur at $t^{\prime}$, then the left- and right-limits agree and $\mathbf{x}_{t^{\prime}-}=\mathbf{x}_{t^{\prime}+}=\mathbf{x}_{t^{\prime}}$. If an event does occur at time $t^{\prime}$ with mark $k$, then the left-limit $\mathbf{x}_{t^{\prime}-}$ can be incremented by $\tilde{\mathbf{E}}\bm{\alpha}_{k}$ to compute $\mathbf{x}_{t^{\prime}+}=\mathbf{x}_{t^{\prime}}$. This increment from left- to right-limit is exact and leverages no discretization assumption.
@@ B.3 Interpretation for Input-Dependent Dynamics
Consider the input-dependent recurrence for an LLH layer, as defined in Eq. 17:
@@ B.4 Forwards and Backwards Zero Order Hold Discretization
In Section 3.3 we highlighted that the ZOH discretization is exact when $\mathbf{u}_{t}$ is held constant over the integration window. This raises a unique design question for DLHPs: what constant value should $\mathbf{u}_{t}$ take on when evolving $\mathbf{x}$ from time $t$ to $t^{\prime}$? For the first layer of the model, the input is zero by construction, so there is no choice to be made—in fact, since $\mathbf{u}$ is constant for the first layer the updates are exact. However, the input is non-zero at deeper layers, and, crucially, varies over the integration period.
We must therefore decide how to select a $\mathbf{u}$ value over the integration period. This should be a value in (or function of) $\{\mathbf{u}_{s}\;|\;s\in[t,t^{\prime})\}$. Note this is because the value at $t^{\prime}$, $\mathbf{u}_{t^{\prime}}$, cannot be incorporated as this would cause a data leakage in our model; while values prior to $t$ would discard the most recent mark. For this work, we explore two natural choices: (i) the input value at the beginning of the interval, $\mathbf{u}_{t}$, and (ii) the left-limit at the end of the interval, $\mathbf{u}_{t^{\prime}-}$. We illustrate the backwards variant in Fig. 2, where in the rightmost panel, we use the $\mathbf{u}_{t^{*}}$ values at each layer, as opposed to $\mathbf{u}_{t_{3}}$. We refer to these options as forwards and backwards ZOH, respectively. All experiments in the main paper utilize backwards ZOH.
It is not obvious a priori which one of these modes is more performant. We therefore conducted an ablation experiment in Table 13. We see that there is little difference between the two methods. We also note that models are learned through this discretization, and so this decision does not mean that a model is “incorrectly discretized” one way or the other, but instead they define subtlety different families of models. Theoretical and empirical investigation of the interpretations of this choice is an interesting area of investigation going forwards, extending the ablations we present in Table 13.
@@ B.5 Theoretical Complexity
We include in Table 4 a brief summary of the theoretical complexity of each of the methods we consider. We break these down by the work, memory complexity and theoretical best parallel application time of the forward pass (used when conditioning on a sequence, the left-hand term of Eq. 2) and evaluating the integral term in Eq. 2 given that the forward pass has been completed (as this is either required by the method, and is nearly always evaluated in conjunction with the forward pass). We then state the limiting best-case theoretical parallelism of the two components.
The reasoning behind this is as as follows:
  • The forward pass of RMTPP, NHP and IFTPP use non-linear RNNs, and hence incur memory and work that is linear in the sequence length, and cannot be parallelized. However, they re-use the computed hidden states to compute the integral term, and hence, while they incur work and memory that scales in the sequence length and number of events, this work can be perfectly parallelized. This results in a best-case parallelism of $\mathcal{O}(L)$ (dominated by the forward pass).
  • SAHP, THP and AttNHP all use self-attention, and hence have a work and memory that scales quadratically in the sequence length, although this work can be parallelized across the sequence length, resulting in logarithmic parallel depth. SAHP and THP re-use embeddings and a parametric decoder, and hence estimating the integral scales like the RNN, and hence the limiting parallelism is still the forward pass. AttNHP is slightly different in that it re-applies the whole independently attention mechanism for each integration point. However, this work is parallelizeable and hence still reduces to a best-case depth of $\mathcal{O}(\log L)$.
  • DLHP is an RNN and hence has linear work and memory in the forward pass, but can be parallelized to a best-case depth of $\mathcal{O}(\log L)$ using the parallel scan. We then re-use the states computed in the forward pass for estimating the integral, which, as with the other RNN methods, is perfectly parallelizable, resulting in a theoretical parallel depth of $\mathcal{O}(\log L)$.
@@
Table 4: Comparison of methods based on memory and compute complexity. We see that our DLHP matches the best performing baseline in all categories. $L$ denotes to the sequence length, and $M$ denotes to the number of Monte Carlo grid points per-event used in evaluating Eq. 2. As IFTPP is an intensity-free method, it does not need to estimate $\int\lambda_{t}\mathrm{d}{}t$ as the other methods do.
@@
RMTPP $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(LM)$ $\mathcal{O}(LM)$ $\mathcal{O}(1)$ $\mathcal{O}(L)$
NHP $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(LM)$ $\mathcal{O}(LM)$ $\mathcal{O}(1)$ $\mathcal{O}(L)$
SAHP $\mathcal{O}(L^{2})$ $\mathcal{O}(L^{2})$ $\mathcal{O}(\log L)$ $\mathcal{O}(LM)$ $\mathcal{O}(LM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log L)$
THP $\mathcal{O}(L^{2})$ $\mathcal{O}(L^{2})$ $\mathcal{O}(\log L)$ $\mathcal{O}(LM)$ $\mathcal{O}(LM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log L)$
AttNHP $\mathcal{O}(L^{2})$ $\mathcal{O}(L^{2})$ $\mathcal{O}(\log L)$ $\mathcal{O}(L^{2}M)$ $\mathcal{O}(L^{2}M)$ $\mathcal{O}(\log L)$ $\mathcal{O}(\log L)$
IFTPP $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(L)$ N/A N/A N/A $\mathcal{O}(L)$
DLHP $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(\log L)$ $\mathcal{O}(LM)$ $\mathcal{O}(LM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log L)$
@@ C.1 Training Details & Hyperparameter Configurations
We apply a grid search for all models on all datasets for hyperparameter tuning. We use a default batch size of 256 for training. For models/datasets that require more memory (e.g., large mark space or long sequences), we reduce the batch size and keep them as consistent as possible among all the models on each dataset. We use the Adam stochastic gradient optimizer (Kingma & Ba, 2015), with a learning rate of 0.01 and a linear warm-up schedule over the first 1% iterations, followed by a cosine decay. Initial experiments showed this setting generally worked well across different models and datasets leads to convergence within 300 epochs. We also clip the gradient norm to have a max norm of 1 for training stability. We use Monte-Carlo samples to estimate the integral in log-likelihood, where we use 10 Monte-Carlo points per event during training.
On the five EasyTPP benchmark datasets and MIMIC-II that are smaller in their scales, we choose an extended grid based on the architecture reported in the EasyTPP paper. Specifically, we search over hidden states size $h$ = {16, 32, 64, 128, 256} for RMTPP, $h$ = {32, 64, 128} for NHP, and $h$ = {16, 32, 64} for IFTPP. For SAHP, THP, and AttNHP, we searched over all combinations of number of $L$ = {1, 2, 3}, hidden state size = {16, 32, 64, 128}, and number of heads = {1, 2, 4}. Finally, for DLHP, we considered combinations for number of layers = {1, 2, 3, 4}, $p$ = {16, 32, 64, 128} and $h$ = {16, 32, 64, 256}. We fixed the activation function as GeLU (Hendrycks & Gimpel, 2016) and apply post norm with layer norm (Ba, 2016). We fix the dropout as 0.1 for DLHP on the five core benchmark datasets, and add dropout = {0, 0.1} to the grid search for the other three datasets. Due to the scale of Last.fm and EHRSHOT datasets, we perform a smaller search over architectures that roughly match the parameter counts for all models at three levels: 25k, 50k, 200k, and choose the model with the best validation results. AttNHP has expensive memory requirements that tends to have smaller batch sizes than other models. We were unable to train any AttNHP on EHRSHOT. The final model architectures used are reported in Table 5(a) and Table 5(b). These configurations are also included in the supplementary code we include.
Table 5: Model architectures for the experiments presented in Table 1
(a) Model architectures for the five EasyTPP benchmark datasets.
@@
IFTPP $h=64$ $h=64$ $h=32$ $h=64$ $h=64$
DLHP $h=64,p=128,l=2$ $h=128,p=128,l=2$ $h=128,p=16,l=4$ $h=32,p=16,l=4$ $h=32,p=32,l=3$
(b) Model architectures for the additional three benchmark datasets.
@@ C.2 Dataset Statistics
We report the statistics of all eight datasets we used in Table 6. We used the HuggingFace version of the five EasyTPP datasets. For all datasets, we further ensure the MTPP modeling assumptions are satisfied that no more than two events occur at the same time (i.e., inter-arrival time is strictly positive), and event times do not lie on grid points that are effectively discrete-time events. Dataset descriptions and pre-processing details are provided in Section C.3.

### v2

@@ B.2 Discretization and Zero Order Hold
The linear recurrence is defined in continuous-time. This mirrors the (M)TPP setting, where event times are not on a fixed intervals. We use the zero-order hold (ZOH) discretization method, to convert the continuous-time linear recurrence into a sequence of closed-form updates, given the integration times, that can also be efficiently computed. We refer the reader to Iserles [2009] for a comprehensive introduction to the ZOH transform.
The main assumption of the ZOH discretization is that the input signal is held constant over the time period being integrated. Under this assumption, it is possible to solve for the dynamics and input matrices that yield the correct state at the end of the integration period. For the LLH dynamics in Eq.˜8, when no events occur in $(t,t^{\prime})$, this becomes
@@
The ZOH does not affect the output or passthrough matrices $\mathbf{C}$ and $\mathbf{D}$. To compute the matrices $\overline{\mathbf{A}}$ and $\overline{\mathbf{AB}}$ however requires computing a matrix exponential and a matrix inverse. Smith et al. [2022] avoid this by diagonalizing the system (also avoiding a dense matrix-matrix multiplication in the parallel scan). The diagonalized dynamics and input matrices are denoted $\bm{\Lambda}$ (a diagonal matrix) and $\bm{\Lambda}\tilde{\mathbf{B}}$ respectively. In this case, Eq.˜16 reduces to
@@
$=(e^{\bm{\Lambda}\Delta t}-\mathbf{I})\tilde{\mathbf{B}}\quad\text{(diagonal matrices commute)}$ (19)
@@
To contextualize, suppose an event occurs at time $t$, Eq.˜19 allows us to exactly (under the constant-input assumption) efficiently evaluate the linear recurrence at subsequent times $t^{\prime}$. We use this extensively in S2P2 to efficiently evaluate the recurrence (and hence the intensity) at the irregularly-spaced event times and times used to compute the integral term.
It should be noted the discretization was done to compute a left-limit $\mathbf{x}_{t^{\prime}-}$ from a previous right-limit $\mathbf{x}_{t}$. Should an event not occur at $t^{\prime}$, then the left- and right-limits agree and $\mathbf{x}_{t^{\prime}-}=\mathbf{x}_{t^{\prime}+}=\mathbf{x}_{t^{\prime}}$. If an event does occur at time $t^{\prime}$ with mark $k$, then the left-limit $\mathbf{x}_{t^{\prime}-}$ can be incremented by $\tilde{\mathbf{E}}\bm{\alpha}_{k}$ to compute $\mathbf{x}_{t^{\prime}+}=\mathbf{x}_{t^{\prime}}$. This increment is exact and leverages no discretization assumption.
@@ B.3 Interpretation for Input-Dependent Dynamics
Consider the input-dependent recurrence for an LLH layer, as defined in Eq.˜13:
@@ B.4 Forwards and Backwards Zero Order Hold Discretization
In Section˜3.2 we highlighted that the ZOH discretization is exact when $\mathbf{u}_{t}$ is held constant over the integration window. This raises a unique design question for S2P2: what constant value should $\mathbf{u}_{t}$ take on when evolving $\mathbf{x}$ from time $t$ to $t^{\prime}$? For the first layer of the model, the input is zero by construction, so there is no choice to be made—in fact, since $\mathbf{u}$ is constant for the first layer the updates are exact. However, the input is non-zero at deeper layers, and, crucially, varies over the integration period.
We must therefore decide how to select a $\mathbf{u}$ value over the integration period. This should be a value in (or function of) $\{\mathbf{u}_{s}\;|\;s\in[t,t^{\prime})\}$. Note this is because the value at $t^{\prime}$, $\mathbf{u}_{t^{\prime}}$, cannot be incorporated as this would cause a data leakage in our model; while values prior to $t$ would discard the most recent event occurrence. For this work, we explore two natural choices: (i) the input value at the beginning of the interval, $\mathbf{u}_{t}$, and (ii) the left-limit at the end of the interval, $\mathbf{u}_{t^{\prime}-}$. Note the end of the interval need not align with an event (crucial for when computing intermediate intensity values). We refer to these options as forwards and backwards ZOH, respectively. We illustrate the backwards variant in Fig.˜2, where in the leftmost panel, we use the $\mathbf{u}_{t-}$ values at each layer to calculate $\mathbf{x}_{t-}$ and $\mathbf{x}_{t}$, as opposed to $\mathbf{u}_{t_{i}}$ for $t_{i}<t<t_{i+1}$. All experiments in the main paper utilize backwards ZOH.
It is not obvious a priori which one of these modes is more performant. We therefore conducted an ablation experiment in Table˜14. We see that there is little difference between the two methods. We also note that models are learned through this discretization, and so this decision does not mean that a model is “incorrectly discretized” one way or the other, but instead they define subtlety different families of models. Theoretical and empirical investigation of the interpretations of this choice is an interesting area of investigation going forwards, extending the ablations we present in Table˜14.
@@ B.5 Theoretical Complexity
We include in Table˜5 a brief summary of the theoretical complexity of each of the methods we consider, broadly grouped by their architectures. We analyze complexity by the work, memory complexity and theoretical best parallel application time of the forward pass (used when conditioning on a sequence, the left-hand term of Eq.˜2) and evaluating the integral term in Eq.˜2 given that the forward pass has been completed (as this is either required by the method, and is nearly always evaluated in conjunction with the forward pass). We then state the limiting best-case theoretical parallelism of the two components.
The reasoning behind the calculated values are as follows:
  • The forward pass of RMTPP, NHP and IFTPP use non-linear RNNs, and hence incur memory and work that is linear in the sequence length, and cannot be parallelized. MHP uses an RNN, but that is logarithmically parallelizable. These models re-use the computed hidden states to compute the integral term, and hence, while they incur work and memory that scales in the sequence length and number of events, this work can be perfectly parallelized. This results in a best-case parallelism of $\mathcal{O}(N)$ (dominated by the forward pass; $\mathcal{O}(\log N)$ for MHP).
  • SAHP, THP and AttNHP all use self-attention, and hence have a work and memory that scales quadratically in the sequence length, although this work can be parallelized across the sequence length, resulting in logarithmic parallel depth. SAHP and THP re-use embeddings and a parametric decoder, and hence estimating the integral scales like the RNN, and hence the limiting parallelism is still the forward pass. AttNHP is slightly different in that it re-applies the whole independently attention mechanism for each integration point. However, this work is parallelizeable and hence still reduces to a best-case depth of $\mathcal{O}(\log N)$.
  • S2P2 is an RNN and hence has linear work and memory in the forward pass, but can be parallelized to a best-case depth of $\mathcal{O}(\log N)$ using the parallel scan. We then re-use the states computed in the forward pass for estimating the integral, which, as with the other RNN methods, is perfectly parallelizable, resulting in a theoretical parallel depth of $\mathcal{O}(\log N)$.
@@
Table 5: Comparison of methods based on memory and compute complexity. We see that our S2P2 matches the best performing baseline in all categories. NN denotes to the sequence length, and $M$ denotes to the number of Monte Carlo grid points per-event used in evaluating Eq.˜2. As IFTPP is an intensity-free method, it does not need to estimate $\int\lambda_{t}\mathrm{d}{}t$ as the other methods do.
@@
RMTPP $\mathcal{O}(N)$ $\mathcal{O}(N)$ $\mathcal{O}(N)$ $\mathcal{O}(NM)$ $\mathcal{O}(NM)$ $\mathcal{O}(1)$ $\mathcal{O}(N)$
NHP $\mathcal{O}(N)$ $\mathcal{O}(N)$ $\mathcal{O}(N)$ $\mathcal{O}(NM)$ $\mathcal{O}(NM)$ $\mathcal{O}(1)$ $\mathcal{O}(N)$
MHP $\mathcal{O}(N)$ $\mathcal{O}(N)$ $\mathcal{O}(\log N)$ $\mathcal{O}(NM)$ $\mathcal{O}(NM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log N)$
IFTPP $\mathcal{O}(N)$ $\mathcal{O}(N)$ $\mathcal{O}(N)$ N/A N/A N/A $\mathcal{O}(N)$
SAHP $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(\log N)$ $\mathcal{O}(NM)$ $\mathcal{O}(NM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log N)$
THP $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(\log N)$ $\mathcal{O}(NM)$ $\mathcal{O}(NM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log N)$
AttNHP $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(\log N)$ $\mathcal{O}(N^{2}M)$ $\mathcal{O}(N^{2}M)$ $\mathcal{O}(\log N)$ $\mathcal{O}(\log N)$
S2P2 $\mathcal{O}(N)$ $\mathcal{O}(N)$ $\mathcal{O}(\log N)$ $\mathcal{O}(NM)$ $\mathcal{O}(NM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log N)$
To validate the scaling properties, we measure the wallclock time for a full forward pass and log-likelihood evaluation on random input sequences with lengths ranging from eight events to over half a million events. The architectures and mark spaces are the same as in the StackOverflow experiments (see Tables˜6(a) and 7).
We note: our EasyTPP PyTorch S2P2 is written in pure PyTorch, and hence is not as optimized as other methods (compared to, for instance, IFTPP, which uses a GPU-optimized implementation of the GRU). We therefore include the runtimes of a standalone JAX S2P2 implementation, which allows for comparable levels of optimization through JIT compilation.
We observe the predicted scaling in practice. NHP scales linearly across all sequence lengths, and is far outpaced by all other methods. The THP scales well before reverting to superlinear scaling, and then runs out of memory. The IFTPP is very fast at shorter runtimes, but quickly reverts to linear scaling, due to its simple but highly optimized implementation and inherently sequential operation. Both S2P2 implementations scale linearly at long sequence lengths, but have near-constant runtime at shorter sequences. At shorter sequence lengths, the more optimized JAX implementation is faster than the unoptimized pure PyTorch implementation. While this indicates that there are additional opportunities to accelerate the PyTorch implementation further (e.g., exploiting kernel fusion or writing a lower-level Triton implementation), these results still confirm that our S2P2 can exploit parallel scans to scale to long sequences more effectively than other methods while retaining strong performance.
Figure 5: Median runtime of various models against increasing sequence lengths when conditioning on a sequence (Algorithm˜1) and for likelihood evaluation (Algorithm˜3) over 20 random seeds (variance negligible). S2P2 is faster across a wide range of sequence lengths. Crosses indicate where THP runs out of memory or IFTPP throws an error.
@@ C.1 Training Details & Hyperparameter Configurations
All baseline models used up-to-date PyTorch implementations, provided by the EasyTPP library [Xue et al., 2023] as of May 2025.
We apply a grid search for all models on all datasets for hyperparameter tuning. We use a default batch size of 256 for training. For models/datasets that require more memory (e.g., large mark space or long sequences), we reduce the batch size and keep them as consistent as possible among all the models on each dataset. We use the Adam stochastic gradient optimizer [Kingma and Ba, 2015], with a learning rate of 0.01 and a linear warm-up schedule over the first 1% iterations, followed by a cosine decay. Initial experiments showed this setting generally worked well across different models and datasets leads to convergence within 300 epochs. We also clip the gradient norm to have a max norm of 1 for training stability. We use Monte-Carlo samples to estimate the integral in log-likelihood, where we use 10 Monte-Carlo points per event during training.
On the five EasyTPP benchmark datasets and MIMIC-II that are smaller in their scales, we choose an extended grid based on the architecture reported in the EasyTPP paper. Specifically, we search over hidden states size $h$ = {16, 32, 64, 128, 256} for RMTPP, $h$ = {32, 64, 128} for NHP, and $h$ = {16, 32, 64} for IFTPP. For SAHP, THP, and AttNHP, we searched over all combinations of number of $L$ = {1, 2, 3}, hidden state size = {16, 32, 64, 128}, and number of heads = {1, 2, 4}. For MHP, we followed their paper to fix $L=4$, then we search over $h$ = {4, 8, 16, 32, 64, 128, 256, 512}. Finally, for S2P2, we considered combinations for number of layers = {1, 2, 3, 4}, $p$ = {16, 32, 64, 128} and $h$ = {16, 32, 64, 256}. We generally found a range of reasonable hyperparameters to yield similar performance on S2P2, while multiple layers were key for performance (intuitively deep stack allows nonlinear and complex dynamics), but there was not a critical dependence on depth.
We fixed the activation function as GeLU [Hendrycks and Gimpel, 2016] and apply post norm with layer norm [Ba, 2016]. We fix the dropout as 0.1 for S2P2 on the five core benchmark datasets, and add dropout = {0, 0.1} to the grid search for the other three datasets. Due to the scale of Last.fm and EHRSHOT datasets, we perform a smaller search over architectures that roughly match the parameter counts for all models at three levels: 25k, 50k, 200k, and choose the model with the best validation results. AttNHP has expensive memory requirements that tends to have smaller batch sizes than other models. We were unable to train any AttNHP on EHRSHOT. The final model architectures used are reported in Tables˜6(a) and 6(b).
Table 6: Model architectures for the five EasyTPP benchmark datasets in Table˜2.
(a) Model architectures for the five EasyTPP benchmark datasets in Table˜2.
@@
S2P2 $h=64,p=128,l=2$ $h=128,p=128,l=2$ $h=128,p=16,l=4$ $h=32,p=16,l=4$ $h=32,p=32,l=3$
(b) Model architectures for the additional three benchmark datasets in Table˜2.
@@ C.2 Dataset Statistics
We report the statistics of all eight datasets we used in Table˜7. We used the HuggingFace version of the five EasyTPP datasets. For all datasets, we further ensure the MTPP modeling assumptions are satisfied that no more than two events occur at the same time (i.e., inter-arrival time is strictly positive), and event times do not lie on grid points that are effectively discrete-time events. Dataset descriptions and pre-processing details are provided in Section˜C.3.